OpenStudy (anonymous):

find f'(x) given f(x)= (sqrt 1+ sqrt 1+x^2)

OpenStudy (anonymous):

\[f(x)=\sqrt 1 + \sqrt {1+x^2}\]??

OpenStudy (anonymous):

ooh maybe \[f(x)=\sqrt{1+\sqrt{1+x^2}}\]

OpenStudy (anonymous):

yes . I know the formula for f ' (x) = lim as z approaches x f(z) - f(x) / z-x correct?

OpenStudy (anonymous):

oh yes

OpenStudy (anonymous):

i wouldn't try that here, it would take all night

OpenStudy (anonymous):

i would use the fact that \[\frac{d}{dx}\sqrt{f(x)}=\frac{f'(x)}{2\sqrt{f(x))}}\]

OpenStudy (anonymous):

ahhh

OpenStudy (anonymous):

put \(f(x)=1+\sqrt{1+x^2}\) and \(f'(x)=\frac{x}{\sqrt{1+x^2}}\) in the above formula \[

OpenStudy (anonymous):

ok I will try that.